1. A pail of water will be swung with a vertical radius of 1.0 m. What will be the minimum speed so that the water inside the pail will not splash out the pail?


1
Expert's answer
2020-11-17T13:03:37-0500

The water inside the pail will not splash out the pail when the centripetal acceleration equals the acceleration due to gravity:


"a_c=g,\\\space\\
\frac{v^2}{R}=g,\\\space\\
v=\sqrt{gR}=3.1\text{ m/s}."